STATE-TO-FIELD VIBRATIONAL ENERGY TRANSFER FROM S_{1} PARA-DIFLUOROBENZENE IN COL-LISIONS WITH SEVERAL INERT GAS PARTNERS.

\underline{T. A. STONE}, D. L. CATLETT, AND C. S. PARMENTER, Indiana University Department of Chemistry, Bloomington, Indiana 47405.

State-to-field vibrational energy transfer (VET) from several levels within the S_{1} (^{1}B_{2u}) manifold of para-Difluorobenzene (pDFB) vapor at 300 K in collisions with a series of gases is currently being examined. Vibrational relaxation by collisions with He, Ne, Ar,^{a} Kr, and Xe from three initial levels with vibrational energies ranging to 1634 cm^{−1} is probed. The resulting fluorescence spectra allow for the determination of absolute rate constants and cross sections for the collisional transfer of vibrational energy into the surrounding S_{1} vibronic field of states.

aD. L. Catlett and C. S. Parmenter, J. Phys. Chem., 95, 2864 (1991).